OpenStudy (anonymous):

Studying for a test and don't remember how to do this. Write the standard equation of the circle with center (-3, 5) and a diameter of 14. (x + 3)2 + (y – 5)2 = 49 (x + 3)2 + (y – 5)2 = 7 (x – 5)2 + (y + 3)2 = 49 (x – 3)2 + (y + 5)2 = 196

OpenStudy (anonymous):

@Hero @dan815 @Nnesha @e.mccormick

Nnesha (nnesha):

\[\huge\rm (x-h)^2+(y-k)^2=r^2\] standard form equation of the circle where (h,k) is the center and r=radius remember radius is half of diameter

OpenStudy (anonymous):

So is it B?

Nnesha (nnesha):

if diameter is 14 then radius would be ?

OpenStudy (anonymous):

7

Nnesha (nnesha):

\[\huge\rm (x-h)^2+(y-k)^2=r^\color{Red}{2}\] yes right so now look at the equation it's r^squared

OpenStudy (anonymous):

So it would be 49

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

So it would be A?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

i think so.:=)

OpenStudy (anonymous):

Thank you so much!!(: